OpenStudy (jade27p):

@iLovePuppiesLol Consider the function f(x) = x2 − 2x + 6. Find the value of f(x) when x = 3 can you please help me

OpenStudy (mitu12):

Plug in 3 for each time you see x

OpenStudy (mitu12):

f(x)=3^2-2(3)+6

OpenStudy (mitu12):

f(3)=9-6+6

OpenStudy (katiebuugcx):

\[3^2-2(3)+6\] \[9-6+6\] \[f(x)=9\]

OpenStudy (mitu12):

Your answer would then be f(3)=9